OpenStudy (anonymous):

2.5 qt/min = ____ gal/h

OpenStudy (tkhunny):

Multiply by 1 as many times as necessary. \(\dfrac{2.5\;qt}{min}\cdot\dfrac{1\;gal}{4\;qt}\cdot What\;Else?\)

OpenStudy (anonymous):

idk

OpenStudy (tkhunny):

Think, ponder. You have quarts. I added a factor that recognizes there are 4 quarts to a gallon. This converts the expression to gallons per minute. What conversion yet remains?

OpenStudy (anonymous):

4 qts

OpenStudy (tkhunny):

No, I already did the quarts to gallons. What units do we want at the end? We need gallons per hour. We had quarts per minutes. I converted to gallons per minute. How do we get hours when we have minutes?

OpenStudy (anonymous):

hour = 60 mins omg lol idk ._. i'm confused :(

OpenStudy (tkhunny):

That is exactly correct. 1 hr = 60 min. What is the value of this expression: \(\dfrac{60\;min}{1\;hr}\)

OpenStudy (anonymous):

idk

OpenStudy (tkhunny):

Did you, or did you not just tell me that 60 min = 1 hour? If I make a ratio out of two things that are equal, what is the result? 5 = 5, so 5/5 = 1 Fred = Fred, so Fred/Fred = 1 Michelle Obama = First Lady, so (Michelle Obama)/(First Lady) = 1 What is the value of \(\dfrac{60\;min}{1\;hr}\)?

OpenStudy (anonymous):

1 -.-